OpenStudy (iwillrektyou):

Help on quadratic equations! Solve 3x^2 + 18x + 15 = 0 by completing the square. 3(x + 3)2 = 6; x = −3 + sqrt2, x = −3 − sqrt2 3(x + 9)2 = 12; x = −11, x = −7 3(x + 3)2 = 12; x = −5, x = −1 3(x + 9)2 = 228; x = −9 + sqrt 76, x = −9 − sqrt 76

OpenStudy (iwillrektyou):

I did it, but I got something wrong. Here is my work:

OpenStudy (misty1212):

HI!!

pooja195 (pooja195):

Find the GCF of all 3 numbers first.

OpenStudy (misty1212):

divide by 3 first

OpenStudy (iwillrektyou):

3x^2 + 18x + 15 = 0 (3x^2 + 18x) + 15 = 0 3(x^2 + 6x) + 15 = 0 3(x^2 + 6x + 9) + 15 = 0 Then, to make it equal: 3(x^2 + 6x + 9) + 15 - 3(9) = 0 3(x^2 + 6x + 9) -12 = 0 3(x+3)^2 - 12 = 0 +12 +12 3(x+3) = 12 Divide 3 on both sides --> x + 3 = 4